OpenStudy (praxer):

Let a,b,c be integers such that gcd(a,b)=1,c>0. Prove that there is an integer x such that gcd(a+bx,c)=1

OpenStudy (amistre64):

proof by example maybe?

OpenStudy (amistre64):

or is this for all c > 0 ?

OpenStudy (amistre64):

euler has an algorithm i believe

OpenStudy (amistre64):

gcd(a,b); assuming a>b gcd(12,5) for example since i need to brush up; 12 and 5 are relatively prime . 12 = 5(2) + 2 5 = 2(2) + 1 <-- this remainder gives the gcd, 2 = 2(1) + 0 is there a simpler way to express that?

OpenStudy (rational):

x = c/gcd(a,c) should work..

OpenStudy (amistre64):

for some a,b ax + by = gcd(a,b) comes to mind as well ... but i cant recall all the thrms at the moment like others :)

OpenStudy (rational):

Let \(x = \dfrac{c}{\gcd(a,c)}\) and consdier the expression \[a+b*\color{blue}{\dfrac{c}{\gcd(a,c)}}\tag{1}\] Let \(d\gt 1\) be a divisor of \(c\) : If \(d\mid a\), then clearly \(d \nmid \color{blue}{\dfrac{c}{\gcd(a,c)}}\) and consequently \(d\nmid (1)\). If \(d\nmid a\), then clearly \(d \mid \color{blue}{\dfrac{c}{\gcd(a,c)}}\) and consequently \(d\nmid (1)\). That means none of the divisors \(\gt 1\) of \(c\) are common to \((1)\). That proves \(\gcd\left(a+b*\color{blue}{\dfrac{c}{\gcd(a,c)}},~~c\right)=1\) for all integers \(a,b,c\)

OpenStudy (praxer):

okay i understood . thank you :)

OpenStudy (rational):

notice that \(\large m+\dfrac{q}{\gcd(m,q)}\) is always coprime to \(\large q\).

OpenStudy (praxer):

okay this is the key concept behind this. Thank you ! :)

OpenStudy (rational):

yw:)

OpenStudy (rational):

there is an error in above proof actually

OpenStudy (rational):

\(\large m+\dfrac{q}{\gcd(m,q)}\) is NOT always coprime to \(\large q\) for example : \(m=2,~q = 2^5\)

OpenStudy (rational):

we need to modify the proof :(

OpenStudy (praxer):

I found a proof in stack exchange just now want to check ???

OpenStudy (rational):

yeah

OpenStudy (rational):

That looks exactly similar to what i am thinking...

OpenStudy (rational):

http://gyazo.com/fedbba3d2e15be1dae4826da85ff3976 looks neat !

OpenStudy (praxer):

but I didnt understand the second part in the proof.

OpenStudy (rational):

thats not second part, that a SECOND PROOF.

OpenStudy (rational):

the snapshot i attached earlier is the complete proof, you're good if that makes sense..

OpenStudy (praxer):

i meant the same. Sorry with the word use :)

OpenStudy (rational):

are you happy with first proof in that mse link ?

OpenStudy (praxer):

how did induction helped in the second proof ??????????? :((